q TFC TVC TC MC AVC ATC
0 $100 $0 $100 -- -- --
1 100 40 140 40 40 140
2 100 60 160 20 30 80
3 100 90 190 30 30 63.33
4 100 124 224 34 31 56
5 100 180 280 56 36 56
6 100 264 364 84 44 60.67
7 100 372 472 108 53.14 67.43
2.1) If the market price is $20, then this firm will maximize profits by producing ________ units of output. (1M)
2.2) If the market price is $84, then this firm will maximize profits by producing ________ unit(s) of output and its profits will be ________. (1M)
2.3) If the market price is $84, then in the long run the firm will _________. (1M)
2.4) If the market price is $34, then in the long run the firm will _________. (1M)
1
Expert's answer
2020-07-27T12:43:03-0400

2.1 Answer: Q = 2 units


Explanation

To maximise profits, firms will produce at the level of output where MR = MC.


In this case, when price = $20, then TR = $20 × Q

= $20Q


"MR = \dfrac {d} {dQ} (TR)"


"= \dfrac {d} {dQ} (20Q)"


"= \$20"


Now, MR = MC = $20 when Q = 2 units.


2.2 Answer: Q = 6 units

π = $140


Explanation

To maximise profits, firms will produce at the level of output where MR = MC.


In this case, when price = $84 then TR = $84 × Q

= $84Q


"MR = \dfrac {d} {dQ} (TR)"


"= \dfrac {d} {dQ} (84Q)"


"= \$84"


Now, MR = MC = $84 when Q = 6 units


for profit : π = TR - TC

TR = $84Q

= $84 × 6

= $504


When Q = 6 units, TC = $364


Therefore, π = $504 - $364

= $140


Thus, when price = $84 the firm makes abnormal profits of $140 at an output level of 6 units.


2.3 Answer: The firm will make normal profits (zero economic profits) in the long run, ceteris paribus.


Explanation

In the long run, supernormal profits will attract new entrants into the industry, assuming existence of weak barriers to entry in the industry. Entrance of rival firms will increase market supply and hence exerting a downward thrust on market price.


The decrease in market price implies a fall in average revenue (AR) and hence a fall in total revenue (TR). These sequence of events will continue until all the supernormal profits are eroded (competed off). The firm will eventually operate at a point where TR = TC and earn normal profits or zero economic profits.


The firm will also reduce its output.


However, if strong barriers to entry exist in the industry, the firm will enjoy monopoly power and will continue to earn supernormal profits in the long run.


2.4 Answer: The firm will continue operating in the long run, and has a chance of earning abnormal profits.


Explanation

From the previous calculations, P = MR.

Therefore, when p = $34, MR = $34.


Since firms maximise profits by producing where MR = MC, then the firm will produce 4 units where MR = MC = $34.


Total profit: π = TR - TC

TR = $34Q

= $34 × 4

= $136

TC = $224 (at 4 units)


Profit (π) = $136 - $224

= -$88

Thus, the firm will earn abnormal losses of $88.


However, average variable costs (AVC) = $31,

and AR > AVC ($34 > $31)


The firm is thus operating above shut down point, and will thus continue operating in the long run. However, if other firms in the industry are also earning abnormal losses, they may exit the industry and hence reducing market supply. Prices will increase resulting in the firm earning abnormal profits in the long run.